# Holmes: Towards Distributed Training Across Clusters with Heterogeneous NIC Environment

### 06 Dec 2023 v1

@@ 2.1 Background and Motivation
2.1.1 RDMA Preliminaries
RDMA is a technology that enables super-efficient data exchange between computers on a network. It allows one computer to access the memory of another without bothering the CPU of either machine. This direct memory access results in lightning-fast data transfers and reduced delay. RDMA substantially improves network performance and efficiency, making it indispensable for applications that demand high-speed data transmission and rapid, low-latency communication. It is particularly crucial in domains such as high-performance computing, cloud computing, and expansive data centers. Two common ways to implement RDMA are InfiniBand and RoCE. InfiniBand is a specialized networking technology, whereas RoCE adapts RDMA to work over Ethernet networks. Both approaches facilitate direct memory access between computers, streamlining data transfer, and communication. The choice between them depends on factors such as network design, specific needs, and the context of the application.
The impact of RDMA technology on LLM training is significant, manifesting in various aspects:
  • Faster Communication Speed: RDMA’s direct memory access accelerates communication operations like model parameter updates and gradient aggregation in distributed environments.
  • More Efficient Data Transfer: RDMA facilitates efficient data transfer, crucial for transmitting parameters and gradients in LLMs. In scenarios with substantial data, efficient transfer enhances training performance.
  • Better Horizontal Scalability: RDMA enhances horizontal scalability, enabling effective LLM training in large GPU clusters without undue concerns about communication bottlenecks.
@@ 2.2 NIC Environments
  • Case 1: Homogeneous Clusters with High-Speed Interconnects. These GPU devices are distributed in clusters that have the same type of RDMA NIC, each equipped with high-speed interconnect (network switch). All GPU devices can communicate through RDMA NICs.
  • Case 2: Heterogeneous Clusters with Absence of High-Speed Interconnects. These GPU devices are distributed across multiple clusters, each equipped with either the same or distinct RDMA NICs, where high-speed interconnects are absent. Communication via RDMA NICs is confined to GPU devices within the same cluster, while GPU devices in different clusters communicate using Ethernet NICs.
@@ 2.3 Problem Description
We describe the core problem addressed in this paper as a scheduling problem for computation tasklets in the heterogeneous NIC environment. The distributed LLM training task is implemented using data and model parallelism techniques, where the entire computational task is decomposed into many parallel groups. Each group represents a set of parallel computations, with each computation viewed as a tasklet. Each tasklet corresponds to the forward and backward computations in a training iteration and is executed on certain GPU devices.
In the heterogeneous NIC environment, the training procedure is constrained by communication. The objective of Holmes is to achieve acceleration by allocating tasklets that require high communication volumes to GPU devices with faster connections. This is achieved by leveraging the characteristics of the netowrk environment, ensuring the maximization of device throughput in the LLM training process. The described method incorporates two parallelism techniques: pipeline and data. Pipeline parallelism involves the concurrent processing of computations in multiple stages. Each device is responsible for activation or gradient computation pertaining to different micro-batches in parallel, and the results can subsequently be communicated or transferred to subsequent stages. On the other hand, data parallelism implies that devices independently compute gradients for various micro-batches, necessitating synchronization of these gradients through communication.
@@
Figure 2: An illustration of parallelism mechanism in Holmes. In this case, a transformer-based model with 6 layers is trained across 2 clusters. Each cluster comprises 2 nodes, each equipped with 4 GPU devices. Communication between Node 1 and Node 2 utilizes InfiniBand, while Node 3 and Node 4 use RoCE. However, there is no high-speed interconnect between the two clusters, and communication between them relies solely on low-speed Ethernet. For parallelism settings, the degrees of data, tensor, and pipeline parallelisms are 2, 2, and 4, respectively. Pipeline parallelism is implemented between the two clusters using Ethernet. The model’s layers are unevenly partitioned into 2 stages and further distributed to different GPU devices across the two clusters. Data parallelism is performed within each cluster using RDMA, and tensor parallelism is implemented within each node using PCI-E.
@@ 2.4 Formalization
  • Let’s take into account a collection of $M$ clusters, which we denote as $\mathbf{C}=\{c_{1},c_{2},\ldots,c_{i},\ldots,c_{M}\}$. It is assumed that each cluster $c_{i}$ comprises $f_{i}$ nodes and that each node is equipped with $G$ devices. The total number of GPU devices is denoted as $N=G\cdot\sum_{i=1}^{M}f_{i}$.
  • We sequentially number clusters, nodes, and GPU devices. In the $i$-th cluster ($0<i\leq M$), the $j$-th GPU device ($0<j\leq G$) in the $k$-th node ($0<k\leq f_{i}$) is denoted as $rank_{G\cdot((\sum_{a=1}^{i-1}f_{a})+k-1)+j}$ in the global context.
@@
  • For pipeline parallelism with a degree of $p$, there are a total of $t\cdot d$ pipeline parallel groups. Each of these groups can be represented as a matrix $[\textbf{PP}]\in\mathbb{Z}_{+}^{(t\ \cdot\ d)\times p}$. Similarly, for tensor and data parallelism with degrees $t$ and $p$, there are a total of $p\cdot d$ tensor parallel groups and $d\cdot t$ data parallel groups, respectively. These two types of group can be represented as matrices $[\textbf{TP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ d)\times t}$ and $[\textbf{DP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ t)\times d}$.
Different scheduling methods may lead to different parallel groups, each with its different communication cost. Our goal is to identify the most efficient scheduling method that minimizes the communication costs of the overall training procedure in the heterogeneous NIC environment.
@@ 3.2 Data Parallelism
We have implemented an auto-select strategy by modifying NCCL [4] and Megatron-LM [3] to support hybrid setups of NICs, including InfiniBand, RoCE, and Ethernet. Our strategy provides functions that create data parallelism groups based on the actual network environment, ensuring that GPU devices within the same data parallelism group are connected to the same type of NIC.
@@ 4.1 Experiment Setup
Hyper-Parameter and Compared Frameworks. The experimental results of Holmes presented in Figure 5, 6, 7, and Table 5 are based on the Self-Adapting Pipeline Partition strategy, with the hyperparameter $\alpha$ set to 1.05. The LLM training frameworks compared with Holmes include Megatron-LM [3], Megatron-LLaMA [2], and Megatron-DeepSpeed [1].
@@ 4.2 Performance Comparison
Figure 6: Performance comparison between Holmes and other mainstream LLM training frameworks: experiment conducted using parameter group 3 on 8 nodes (4 nodes equipped with RoCE NICs and 4 nodes equipped with IB NICs)
@@ 4.4 Discussion
In this section, we undertake comprehensive experiments from several aspects to address the questions posed at the beginning of this section. Firstly, we evaluate the performance of Holmes in various NIC environments, constituting a pivotal experiment within this paper. Our findings demonstrate that Holmes attains a performance level that competes favorably in heterogeneous NIC environments, closely approaching the efficiency achievable in homogeneous RDMA-capable networks (InfiniBand or RoCE), and surpassing the training efficiency observed in purely Ethernet environments. Subsequently, we conduct an in-depth comparative analysis of the different components of Holmes, substantiating their effectiveness. Finally, we compare Holmes with other mainstream LLM training frameworks, thereby verifying its advancements.
@@ 5 Related Work
5.1 Large Language Models
Large language models (LLMs) have attracted considerable attention in recent years, marking a substantial advancement in the field of NLP. Their origins can be traced back to the introduction of the Transformer architecture [36], which, relying on self-attention mechanisms, brought about a revolution in deep learning, particularly in the realm of understanding and generation of natural languages. Notable examples of LLMs include the GPT series [22] and BERT [8]. The fundamental concept underlying LLMs is pre-training on extensive text corpora, which equips the models with a broad grasp of language, encompassing syntax, semantics, and world knowledge. Subsequently, these models undergo fine-tuning for specific tasks, leading to state-of-the-art performance in a series of NLP tasks like language modeling, text classification, machine translation, and question-answering.
@@ 5.2 Distributed Training Frameworks
Distributed training frameworks are essential for efficient and scalable Large Language Model (LLM) training. They enable parallel task execution across multiple devices, addressing computational and memory demands of large models. LLMs are usually trained in data centers with homogeneous high-speed RDMA NICs connecting GPUs. However, existing frameworks often neglect training LLMs in heterogeneous NIC environments.
Some popular distributed training frameworks tailored for Large Language Models (LLMs) include Megatron-LM [32, 25, 3], Megatron-LLaMA [2], and DeepSpeed [27, 30, 28, 1]. Megatron-LM, developed by NVIDIA, efficiently handles training of very large models using a combination of model parallelism and data parallelism across multiple GPUs and nodes. This approach enables LLM training by sharing model parameters among GPUs and processing data in parallel [25, 39]. Megatron-LM excels in minimizing synchronization delays during training through optimized communication between nodes and GPUs, utilizing efficient communication and synchronization techniques. Megatron-LLaMA, an extension of Megatron-LM, enhances the framework by integrating a standardized LLaMA implementation and introducing an efficient communication-computation parallelism method. Another notable framework is DeepSpeed, an open-source library from Microsoft designed for efficient training of large-scale deep learning models. DeepSpeed features memory-efficient training, accommodating models larger than GPU memory capacity through techniques like mixed-precision training, gradient accumulation, and offloading optimizer states. Its ZeRO optimization technology further improves memory efficiency by distributing model weights and optimizer states among GPUs and nodes, reducing memory usage and enabling training of even larger models [27, 30].
@@ 6 Conclusion
In this paper, we introduce Holmes, a meticulously designed framework designed for training LLM across multiple GPU clusters. Our empirical studies indicate that the proposed training framework can achieve performance levels in the heterogeneous NIC environment comparable to those possible with homogeneous RDMA NICs. Holmes is considered a significant advancement in making LLM training more accessible, thereby expanding the potential for efficient scaling within the broader research community.

### 07 Dec 2023 v2

@@ 2.1 Background and Motivation
2.1.1 Communication Technologies
In addition to Ethernet, which is a widely used protocol for network connection, RDMA, NVLink, and PCI-E are three distinct technologies or protocols used to support varying degrees of data transmission and communication [9, 17, 10].
  • RDMA is a network technology enabling direct memory access between computer systems without CPU involvement in data transfer. It achieves high-performance remote memory access by bypassing the CPU, reducing latency.
  • NVLink is a high-speed technology connecting multiple GPUs for enhanced parallel computing. It provides a high-bandwidth, low-latency communication channel with direct memory access, improving data transfer efficiency in multi-GPU systems.
  • PCI-E is a high-speed serial bus standard connecting diverse hardware in computers, including graphics cards, storage, and network devices. It offers a versatile, high-bandwidth, and low-latency data transfer channel through point-to-point connections. While PCI-E is commonly used for various hardware, it may have limitations in data transfer efficiency compared to technologies like NVLink, specifically designed for GPU-to-GPU communication..
In summary, RDMA enables high-performance network communication, NVLink enhances communication between GPUs, and PCI-E is a versatile and widely used internal computer bus standard. These technologies are crucial in various domains and applications, particularly for distributed training.
Figure 2: An illustration of the scheduling problem in heterogeneous NIC environments, with the objective of minimizing communication costs for distributed LLM training tasks.
@@ 2.2 NIC Environments
  • Case 1: Homogeneous Clusters with High-Speed Interconnects. These GPU devices are distributed among multiple clusters, each equipped with the same type of RDMA NIC and connected through high-speed interconnects. All GPU devices can communicate through RDMA NICs.
  • Case 2: Heterogeneous Clusters with Absence of High-Speed Interconnects. These GPU devices are distributed among multiple clusters, each equipped with either the same or distinct types of RDMA NIC. Notably, there are no high-speed interconnects between these clusters. Communication via RDMA NICs is limited to GPU devices within the same cluster, whereas GPU devices located in different clusters communicate using Ethernet NICs.
@@ 2.3 Problem Description
As illustrated in Figure 2, this paper focuses on scheduling computation tasklets in the heterogeneous NIC environment with varying bandwidth values. The distributed LLM training task is implemented using data and model parallelism techniques, where the entire computational task is decomposed into many parallel groups. Each group represents a set of parallel computations, with each computation viewed as a tasklet. Each tasklet corresponds to the forward and backward computations in a training iteration and is executed on certain GPU devices.
In the heterogeneous NIC environment, the training procedure is constrained by communication. The objective of Holmes is to achieve acceleration by allocating tasklets that require high communication volumes to GPU devices with faster connections. This is achieved by leveraging the characteristics of the network environment, ensuring the maximization of device throughput in the LLM training process. The described method incorporates two parallelism techniques: pipeline and data. Pipeline parallelism involves the concurrent processing of computations in multiple stages. Each device is responsible for activation or gradient computation pertaining to different micro-batches in parallel, and the results can subsequently be communicated or transferred to subsequent stages. In addition, data parallelism implies that devices independently compute gradients for various micro-batches, necessitating synchronization of these gradients through communication.
@@ 2.4 Formalization
  • Let’s take into account a collection of $M$ clusters, which we denote as $\mathbf{C}=\{c_{1},c_{2},\ldots,c_{i},\ldots,c_{M}\}$. It is assumed that each cluster $c_{i}$ comprises $f_{i}$ nodes and that each node is equipped with a constant number of $G$ devices. The total number of GPU devices is denoted as $N=G\cdot\sum_{i=1}^{M}f_{i}$.
  • We sequentially number clusters, nodes, and GPU devices. In the $i$-th cluster ($0<i\leq M$), the $j$-th GPU device ($0<j\leq G$) in the $k$-th node ($0<k\leq f_{i}$) is denoted as $\operatorname{rank}_{G\cdot((\sum_{a=1}^{i-1}f_{a})+k-1)+j}$ in the global context.
@@
  • For pipeline parallelism with a degree of $p$, there are a total of $t\cdot d$ pipeline parallel groups. Each of these groups can be represented as a matrix $[\textbf{PP}]\in\mathbb{Z}_{+}^{(t\ \cdot\ d)\times p}$. Similarly, for tensor and data parallelism with degrees $t$ and $p$, there are a total of $p\cdot d$ tensor parallel groups and $p\cdot t$ data parallel groups, respectively. These two types of group can be represented as matrices $[\textbf{TP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ d)\times t}$ and $[\textbf{DP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ t)\times d}$.
Different scheduling methods may lead to different parallel groups, each with its different communication costs. Our goal is to identify the most efficient scheduling method that minimizes the communication costs of the overall training procedure in the heterogeneous NIC environment.
@@ 3 Holmes Design
Figure 3: An illustration of parallelism mechanism in Holmes. In this case, a transformer-based model with 6 layers is trained across 2 clusters. Each cluster comprises 2 nodes, each equipped with 4 GPU devices. Communication between Node 1 and Node 2 utilizes InfiniBand, while Node 3 and Node 4 use RoCE. However, there is no high-speed interconnect between the two clusters, and communication between them relies solely on low-speed Ethernet. For parallelism settings, the degrees of data, tensor, and pipeline parallelisms are 2, 2, and 4, respectively. Pipeline parallelism is implemented between the two clusters using Ethernet. The model’s layers are unevenly partitioned into 2 stages and further distributed to different GPU devices across the two clusters. Data parallelism is performed within each cluster using RDMA, and tensor parallelism is implemented within each node using NVLink.
@@ 3.2 Data Parallelism
We have implemented an auto-select strategy by modifying NCCL [4] and Megatron-LM [3] to support hybrid setups of NICs, including InfiniBand, RoCE, and Ethernet. As shown in Figure 4, our strategy provides functions that create data parallelism groups based on the actual network environment. It establishes independent communication channels for each group of devices, ensuring that GPU devices within the same data parallelism group are connected to the same type of NIC.
@@
Figure 4: Code snippet of creating data parallel groups across clusters within the heterogeneous NIC environment.
@@ 4.1 Experiment Setup
Hyper-Parameter and Compared Frameworks. The experimental results of Holmes presented in Figure 7, 8, 9, and Table 5 are based on the Self-Adapting Pipeline Partition strategy, with the hyperparameter $\alpha$ set to 1.05. The LLM training frameworks compared with Holmes include Megatron-LM [3], Megatron-LLaMA [2], and Megatron-DeepSpeed [1].
@@ 4.2 Performance Comparison
Figure 8: Performance comparison between Holmes and other mainstream LLM training frameworks: experiment conducted using parameter group 3 on 8 nodes (4 nodes equipped with RoCE NICs and 4 nodes equipped with IB NICs).
@@ 4.4 Discussion
In this section, we conduct comprehensive experiments to address the initial questions. Firstly, we evaluate Holmes’s performance across various NIC environments, a key experiment in this paper. Our findings show that Holmes competes favorably in heterogeneous NIC environments, approaching the efficiency of homogeneous RDMA-capable networks (InfiniBand or RoCE) and surpassing training efficiency in purely Ethernet environments. Next, we perform a detailed comparative analysis of Holmes’s components, substantiating their effectiveness. Finally, we compare Holmes with other mainstream LLM training frameworks to validate its advancements.
@@ 5.1 Distributed Training Frameworks
Distributed training frameworks are essential for efficient and scalable LLM training. They enable parallel task execution across multiple devices, addressing computational and memory demands of large models. LLMs are usually trained in data centers with homogeneous high-speed RDMA NICs connecting GPUs. However, existing frameworks often neglect training LLMs in heterogeneous NIC environments.
Some popular distributed training frameworks tailored for LLMs include Megatron-LM [30, 23, 3], Megatron-LLaMA [2], and DeepSpeed [25, 28, 26, 1]. Megatron-LM, developed by NVIDIA, efficiently handles training of very large models using a combination of model parallelism and data parallelism across multiple GPUs and nodes. This approach enables LLM training by sharing model parameters among GPUs and processing data in parallel [23, 36]. Megatron-LM excels in minimizing synchronization delays during training through optimized communication between nodes and GPUs, utilizing efficient communication and synchronization techniques. Megatron-LLaMA, an extension of Megatron-LM, enhances the framework by integrating a standardized LLaMA implementation and introducing an efficient communication-computation parallelism method. Another notable framework is DeepSpeed, an open-source library from Microsoft designed for efficient training of large-scale deep learning models. DeepSpeed features memory-efficient training, accommodating models larger than GPU memory capacity through techniques like mixed-precision training, gradient accumulation, and offloading optimizer states. Its ZeRO optimization technology further improves memory efficiency by distributing model weights and optimizer states among GPUs and nodes, reducing memory usage and enabling training of even larger models [25, 28].
@@ 6 Conclusion
In this paper, we present Holmes, a carefully crafted framework for LLM training across multiple GPU clusters. Our empirical studies show that this framework performs comparably to homogeneous RDMA NICs in a heterogeneous NIC environment. Holmes represents a notable advancement, enhancing accessibility to LLM training and facilitating efficient scaling within the broader research community.